Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report 2019-08-28

Night Report
Observers: Emma Callis, Seán Brennan
Support Team: Ihanec, Gromadzki
TAT: Maria Teresa Botticella, Hanin Kuncarayakti, Janet Chen

Calibrations: 
EFOSC:
Daytime - 
    pessto_bias.obx    
    pessto_calib.obx
Evening Twilight - 
    Skyflats UBz, ran out of twillight, will try get VRi again

Weather: 
Start:  Clear, seeing 0.7"
Middle: Clear, seeing 0.7" - 2"
End:  Clear, seeing 1"


Comments:
Mainly using Las Campanas DIMM 

Big issue with the telescope, TO saying azimuth went to standby, and when trying to initialise system there was something wrong with the encoders, 
telescope wasn't reaching the initialisation point
Couldn't resolve issue tonight, will have to wait for daytime technical team


EFOSC
 Time (UT)    
 Target     Observation Seeing     Comments
 23:20Image Analysis   0.66 
 23:32 SN2019ltw Flw: Gr13 1x900 0.73 
 23:54 AT2019otb Cls: Gr13 1x300 0.7 Blue featureless
 00:12 AT2019ojd Cls: Gr13 1x900  0.77 Double checked target location with EFOSCFASTPHOT, SNID=IIP z~0.046
 00:32 STND EG131 Gr11 + Gr13 + Gr16 0.91 
 00:58 AT2019omk Cls: Gr13 1x300 0.8  SNID=Ia z~0.038
 01:17 AT2019oqq Cls: Gr13 1x1500 1.12 Not visible in Acq. Image with EFOSCFASTPHOT
 01:23 AT2019onu Cls: Gr13 1x900 0.9 Not visible in Acq. Image with EFOSCFASTPHOT
 01:28 AT2019oop Cls: Gr13 1x900 1.08 Blue featureless
 01:53 AT2019ovn Cls Gr13 1x1500 1.15 Not visible in Acq. Image
 01:58 AT2019oto Cls: Gr13 1x900 1.13 SNID=Ia z~0.1 ?
 02:20 AT2019oje Cls: Gr13 1x900 1.3 SNID=IIb z~0.04
 02:44 SN2019odp Flw: Gr11 1x1800   0.86 Accidentally called it SN20190dp in header for both Gr11 spec
 03:19 SN2019odp Flw: Gr11 1x1800  1.22 
 03:55 SN2019odp Flw: Gr16 1x1800  0.94 
 04:41 SN2019odp Flw: Gr16 1x1800 1.19 
 05:25 AT2019nwl Cls: Gr13 1x900 1.37 SNID=Ia z~0.8
 05:47 SN2019muj Flw: Gr11 1x2700 1.13 
 06:37 SN2019muj Flw: Gr11 1x2700 1.1 Aborted @07:10 due to technical issue
 07:10 ------ --- Technical issue, azimuth went to standby, problem with encoders, TO went to   telescope
 07:58 --- ---  TO had to reboot system
 08:24    Unable to resolve issue tonight, night cut short




Comments